a binary compound consist of z and m elements, weight of z is 14 and that of m is 42. if 0.1 mole of the compound is decomposed completely, the total number of atoms obtained is the same as the number of atom in 12.8 grams of oxygen gas. the amount of m obtained in grams upon decomposition is exactly 9 times the amount of z. what is the molecular formula of the compound
1
Expert's answer
2020-08-19T04:32:17-0400

n(O2)=m/M=12.8/32=0.4 (mol)

n(O)=0.4×2=0.8 (mol).

It is 8 times bigger than 0.1 mol, so from 1 molecule of compound we obtain 8 atoms.

If we assume that the formula of a compound is zxmy, we may say that

x+y=8

or y=8-x.

We may also say that masses of elements in the compound depend on their atomic weight and number of atoms:

for z it is equal to 14x,

for m it is equal to 42y.

Taking into account that mass of m is equal to 9 masses of z we obtain the second equation:

9×14x = 42y

or

126x=42×(8-x),

126x=336-42x,

168x=336,

x=2,

y=8-2=6.

Answer: z2m6.